OpenStudy (anonymous):

2. Which value for x makes the sentence true? 4x - 2 = 18 x = 24 x = 4 x = 20 x = 5

OpenStudy (black_ninja123):

4*5=20 20-2=18 x=5

OpenStudy (anonymous):

thanks :D

OpenStudy (black_ninja123):

Your welcome

OpenStudy (lizz123):

4x - 2 = 18 So the opposite of subtraction is addition so you have to add 2 on both sides. 4x - 2 = 18 4x + 2 = + 2 4x _______ = ______ -2+2 is cancelled and add 18 +2 which equals 20. So your new equation would be: 4x = 20 Now you have to divide by 4. 4x = 20 _____ _____ 4 4 Since 4x divide that by isn't a way to divide it is cancelled and now you have 20 divide it by 4. which you answer will be 5. Therefore, x=5.